Drain Water Disposal Company has a long-term contract with several large cities to collect garbage and trash from residential customers. To facilitate the collection, Drain places a large plastic container with each household. Several years ago, the company decided to manufacture its own containers as a cost-saving measure. To help ensure cost efficiency, a standard cost system was Installed in the plant. The following standards have been established for the product's variable inputs.

DURATION EQUIPMENT ALLOWED : CALCULATOR CHARACTERISTICS :120 MINUTES (2 HOURS) CLOSED BOOK L Standard Costing (30%) Drain Water Disposal Company has a long-term contract with several large cities to collect garbage and trash from residential customers. To facilitate the collection, Drain places a large plastic container with each household. Several years ago, the company decided to manufacture its own containers as a cost-saving measure. To help ensure cost efficiency, a standard cost system was Installed in the plant. The following standards have been established for the products variable inputs Direct materials Direct labor Variable overhead Fixed overhead Standard usage 12 lbs 1.70 hours 1.70 hours 1.70 hours Standard price 3.50/lb. $11.00/hour $ 3.00/hour $ 2.00/hour Standard cost $42.00 $ 18.70 $ 5.10 $ 3.40 During the first week, the company had the following actual results: 6,000 units Units produced Actual direct labor costs Actual direct labor hours Direct materials used in production Actual direct materials price Actual variable overhead costs Actual fixed overhead costs 118,800 10,800 hours 69,000 lbs. $3.55/lb. $29,700 $27,000 Required 1. Compute direct materials price, quantity, and total variances 2. Compute direct labor price, quantity, and total variances 3. Compute the total overhead variance
Answer #1

1. Material Variance:

a.

(1) (2) (3)
SQSP AQSP AQAP

(6,000*12) * 3.5

=252,000

69,000* 3.5

= 241,500

69000* 3.55

=244,950

SQ= 6000*12 = 72,000

(i). Material Price Variance (2-3) = 241,500 - 244,950 = $3,450 (A)

(ii). Material Usuage Variance(1-2) = 252,000 - 241,500 = $10,500 (F)

(iii). Material Cost Variance (1-3) = 252,000 - 244,950 = $7050 (F)

2. Labour Variance:

  (1) (2) (3)
  SRSH SRAH ARAH

(6000*1.70) * 11

=10,200*11 = 112,200

10,800* 11

=118,800

$118,800

(i). Direct Labor Price (2-3) = 118,800 - 118,800 = 0

(ii). Direct Quantity Variance (1-2) = 112,200 - 118,800 = $6,600 (A)

(iii). Direct Cost Variance (1-3) = 112,200 - 118,800 = $6,600 (A)

3. Total overhead Variance -

Total Actual Cost - Total Standard Cost

= Total Actual Cost = Total Actual Fixed Overhead + Total Actual Variable Overhead  

= $29,700 + $27,000 = $56,700

Total Standard Cost = Total Applied Variable Overhead + Total Applied Fixed Overhead

= $30,600 + $20,400 = $51,000

Total Overhead Variance = $56,700 - $51,000 = $5,700